The perimiter of a rectangle is 30m.
It's area is 56m2.
What is its length?

Answer by ankor@dixie-net.com(15652)

The perimeter of a rectangle is 30m.
2L + 2W = 30
Simplify, divide by 2
L + W = 15
W = (15-L)
:
It's area is 56m2.
L*W = 56
Replace W with (15-L)
L(15-L) = 56
15L - L^2 = 56
Form a quadratic equation on the right
0 = L^2 - 15L + 56
Factors to
(L-7)(L-8) = 56
Two solutions
L = 7; then W = 8
L = 8; then W = 7